You are hired by Murphey Farms to decrease the backfat in their hogs. The current hog population has a mean backfat content of 17.8mm. You selected the hogs with the smallest amount of backfat, mated them and obtained the progeny generation. The selected hogs had a mean backfat content of 11.8mm. You know that the heritability for percent backfat in hogs is 0.53.

What is the selection differential? Round your values properly to 2 decimal digits.

What is the expected response to selection? Round your values properly to 2 decimal digits.

What is the expected mean of the progeny generation? Round your values properly to 2 decimal digits.
Answer #1

Answer 1.

Selection differential is difference between mean of selected parents (Ts) and actual mean, T (with all the parents).

S = Ts-T

= (11.8 - 17.8) = 6 mm

Answer 2.

Expected response to selection = R = (T' - T) = h2 * S; where h = heritability and S = 6mm (calculated in q.1)

R = (0.53 * 6) = 3.18 mm

Answer 3.

Expected mean of progeny generation = T'

R = (T' - T); where, T= 17.8mm and R = 3.18mm

T' = (3.18 + 17.8) = 20.98 mm